OpenStudy (anonymous):

Solve for x in (x/3)-(1/3)=(2/x)

OpenStudy (turingtest):

what happened to our other question? you just sort of left like you gave up...

OpenStudy (anonymous):

I posted on there that i got the answer which was -1/6. I tried to find the least common multiple which is 6 but when i went to work it, i got lost. it would be 6(x/3-1/3=2/x) right?

OpenStudy (turingtest):

not quite, look at the denominators: 3 3 x ^^^ those are our denominators we don't need to repeat numbers to find the LCM so which are the only two terms which are different? 3 and x so the LCM is 3x

OpenStudy (anonymous):

Oh okay hold up let me work it.

OpenStudy (anonymous):

Forget it idc its too hard.

OpenStudy (turingtest):

can you do something like simplify \[3x(\frac x3)\]?